Please help me solve this linear equation using substitution
0.6x - 0.2y = -0.2
-0.03x - 0.07y = 0.17
Thank you.

Answer by DrBeeee(684)

Given:
(1) .6x -.2y = -.2
(2) -.03x -.07y = .17
First, make it easy on yourself by getting rid of the decimals. Multiply (1) by 10 and (2) by 100 and get
(3) 6x - 2y = -2
(4) -3x - 7y = 17
Now solve (3) for y
(5) 6x + 2 = 2y or
(6) y = 3x + 1
Now substitute (6) into (4) and get
(7) -3x - 7(3x+1) = 17 or
(8) -3x -21x -7 = 17 or
(9) -24x = 24 or
(10) x = -1
Now use (6) to find y
(11) y = 3(-1) + 1 or
(12) y = -2
Check this pair using (1).
Is (.6(-1) - .2(-2) = -.2)?
Is (-.6 + .4 = -.2)?
Is (-.2 = -.2)? Yes
Answer: The solution pair is (-1,-2)